# How Good Is It? Evaluating the Efficacy of Common versus Domain-Specific Prompts on Foundational Large Language Models

### v1

@@ Abstract
Recently, large language models (LLMs) have expanded into various domains. However, there remains a need to evaluate how these models perform when prompted with commonplace queries compared to domain-specific queries, which may be useful for benchmarking prior to fine-tuning domain-specific downstream tasks. This study evaluates LLMs, specifically Gemma-2B and Gemma-7B, across diverse domains, including cybersecurity, medicine, and finance, compared to common knowledge queries. This study employs a comprehensive methodology to evaluate foundational models, encompassing problem formulation, data analysis, and the development of novel outlier detection techniques. This methodological rigor enhances the credibility of the presented evaluation frameworks. This study focused on assessing inference time, response length, throughput, quality, and resource utilization and investigated the correlations between these factors. The results indicate that model size and types of prompts used for inference significantly influenced response length and quality. In addition, common prompts, which include various types of queries, generate diverse and inconsistent responses at irregular intervals. In contrast, domain-specific prompts consistently generate concise responses within a reasonable time. Overall, this study underscores the need for comprehensive evaluation frameworks to enhance the reliability of benchmarking procedures in multidomain AI research.
@@ V Conclusion
This study investigates the inference behavior of foundation models of varying sizes under common and domain-specific prompts, such as those related to cybersecurity, medical, and finance domains. This study examines these behaviors under conditions in which response lengths are both restricted and unrestricted. We present a framework to assess large foundational models in terms of domain understanding and outlier formulation. The results indicate that model size and types of prompts used for inference significantly influence response length and quality, as larger datasets for training provide more information across various domains. In addition, common prompts, which include different types of queries, generate diverse responses and may result in inconsistent response lengths when the same prompt is used multiple times or in different ways. In contrast, domain-specific prompts consistently generate concise responses. Therefore, we recommend eliminating irrelevant domains in the language model information prior to fine-tuning domain-specific tasks.

### v2

@@ Abstract
Recently, large language models (LLMs) have expanded into various domains. However, there remains a need to evaluate how these models perform when prompted with commonplace queries compared to domain-specific queries, which may be useful for benchmarking prior to fine-tuning for domain-specific downstream tasks. This study evaluates LLMs, specifically Gemma-2B and Gemma-7B, across diverse domains, including cybersecurity, medicine, and finance, compared to common knowledge queries. This study utilizes a comprehensive methodology to assess foundational models, which includes problem formulation, data analysis, and the development of ThroughCut, a novel outlier detection technique that automatically identifies response throughput outliers based on their conciseness. This methodological rigor enhances the credibility of the presented evaluation frameworks. This study focused on assessing inference time, response length, throughput, quality, and resource utilization and investigated the correlations between these factors. The results indicate that model size and types of prompts used for inference significantly influenced response length and quality. In addition, common prompts, which include various types of queries, generate diverse and inconsistent responses at irregular intervals. In contrast, domain-specific prompts consistently generate concise responses within a reasonable time. Overall, this study underscores the need for comprehensive evaluation frameworks to enhance the reliability of benchmarking procedures in multidomain AI research.
@@ I Introduction
  4. 4.
    We propose a novel outlier detection technique, termed ThroughCut, which automatically identifies response throughput outliers by assessing their conciseness.
@@ V Conclusion
This study investigates the inference behavior of foundation models of varying sizes under common and domain-specific prompts, such as those related to cybersecurity, medical, and finance domains. This study examines these behaviors under conditions in which response lengths are both restricted and unrestricted. We present a framework to assess large foundational models in terms of domain understanding and outlier formulation. The results indicate that model size and types of prompts used for inference significantly influence response length and quality, as larger datasets for training provide more information across various domains. In addition, common prompts, which include different types of queries, generate diverse responses and may result in inconsistent response lengths when the same prompt is used multiple times or in different ways. In contrast, domain-specific prompts consistently generate concise responses. Therefore, we recommend eliminating irrelevant domains in the language model information prior to fine-tuning domain-specific tasks. For example, when provided with sufficient datasets to fine-tune the 2 billion parameter model for a cybersecurity downstream task, we advocate for a full fine-tuning approach rather than employing a parameter-efficient technique. This method involves the elimination of irrelevant domains, allowing the target domain to become predominant. Such an approach preserves the model’s size while facilitating the generation of concise and consistent responses in a minimal amount of time.